OpenStudy (anonymous):

Free medal given!! HELP! solve each equation on the interval [0,2pi) sin3x=-1/2

OpenStudy (anonymous):

@bibby

OpenStudy (anonymous):

@misty1212

OpenStudy (anonymous):

2cos^2 x + 5cos x + 2 = (2cos x + 1)(cos x + 2) = 0 cos x = -1/2 x = 2pi/3, or 4pi/3

OpenStudy (anonymous):

i'm talking about for sine

OpenStudy (anonymous):

@nikato

OpenStudy (freckles):

so you know how to solve sin(u)=-1/2 use the unit circle

OpenStudy (freckles):

by the way if x is between 0 and 2pi so if u=3x then u/3=x so u/3 is between 0 and 2 pi therefore u is between 0(3) and 2pi(3) or if we simplify that u is between 0 and 6pi so solve sin(u)=-1/2 on the interval [0,6pi] first then we will solve for x

OpenStudy (anonymous):

x will be 4pi/3 and 5pi/3

OpenStudy (freckles):

what did you solve?

OpenStudy (freckles):

are you solving this for u first: sin(u)=-1/2 on [0,6pi]?

OpenStudy (anonymous):

i dont understand how to do that. i'm just diong what i know. the question is sin3x=-1/2 i want to get the steps in the fastest quickest way popssible

OpenStudy (freckles):

so have you ever used the unit circle? because that is normally how you solve something like sin(u)=-1/2 on the interval I gave

OpenStudy (freckles):

you find for what angles you have the y value is -1/2

OpenStudy (anonymous):

i have but im not really sure. can you show me the steps you will take

OpenStudy (freckles):

for example I will find one value for you \[\sin(u)=\frac{-1}{2} \\ \text{ when } u=\frac{7 \pi}{6}\] can you find any other values where the angle gives you the y value is -1/2

OpenStudy (freckles):

all you are doing is using a table

OpenStudy (freckles):

look at the unit circle and just observe when the y-coordinate is -1/2

OpenStudy (freckles):

hint: look all in the 4th quadrant

OpenStudy (freckles):

also*

OpenStudy (anonymous):

11pi/6

OpenStudy (freckles):

yes

OpenStudy (freckles):

there are some more values for u since we are looking from u=0 to u=6pi

OpenStudy (anonymous):

no it stops at 2pi

OpenStudy (freckles):

\[u=\frac{7\pi}{6}, 2\pi+\frac{7\pi}{6},4\pi+\frac{7\pi}{6} \\ u=\frac{11\pi}{6},2\pi+\frac{11\pi}{6},4\pi+\frac{11\pi}{6}\] now you can finally solve for x replace u with 3x

OpenStudy (freckles):

and solve those 6 equations

OpenStudy (freckles):

as i said earlier x is between 0 and 2pi but we replaced 3x with u which means u=3x and x=u/3 so u/3 is between 0 and 2pi so u is actually between 0 and 6pi so we had to solve sin(u)=-1/2 on the interval [0,6pi]

OpenStudy (anonymous):

i dont understand can yuou explain it to me in equation form?

OpenStudy (freckles):

do you know how to replace u with 3x?

OpenStudy (anonymous):

i would go about this by doing 3x=4pi/3+2npi

OpenStudy (freckles):

\[3x=\frac{7\pi}{6},2\pi+\frac{7\pi}{6},4\pi+\frac{7\pi}{6} \\ 3x=\frac{11\pi}{6},2\pi+\frac{11\pi}{6},4\pi+\frac{11\pi}{6}\]

OpenStudy (freckles):

now solve for x

OpenStudy (anonymous):

the explain it ot me b making equations first sin is negative in the 3rd and 4th quariants and isnt that 4pi3 and 5pi/3

OpenStudy (freckles):

are you chaging the equation to sin(x)=-sqrt(3)/2 ?

OpenStudy (anonymous):

no

OpenStudy (freckles):

then how are you getting those values

OpenStudy (anonymous):

i don tknow if its right. i was asking you

OpenStudy (freckles):

so you want to resolve sin(u)=-1/2 then?

OpenStudy (freckles):

again we are looking for angle u such that the y-coordinate is -1/2

OpenStudy (freckles):

we discovered the values 7pi/6 and 11pi/6 satisfy this

OpenStudy (freckles):

then we had u was between 0 and 6pi so that means we have three rotation of answers for each

OpenStudy (freckles):

so we have 6 answers in all

OpenStudy (freckles):

which were written above

OpenStudy (freckles):

but now you have to solve for x

OpenStudy (freckles):

by dividing 3 on both sides

OpenStudy (freckles):

\[x=\frac{7\pi}{6(3))},\frac{2\pi}{3}+\frac{7\pi}{6(3)},\frac{4\pi}{3}+\frac{7\pi}{6(3)} \\ x=\frac{11\pi}{6(3)},\frac{2\pi}{3}+\frac{11\pi}{6(3)},\frac{4\pi}{3}+\frac{11\pi}{6(3)} \\ \]

OpenStudy (freckles):

I think you still don't get it. I will try to explain one more time. sin(3x)=-1/2 where 0<=x<=2pi Let 3x=u. We are going to put everything in terms of u including the inequality. sin(u)=-1/2 where 0<=u/3<=2pi since u=3x => u/3=x solving the inequality gives 0<=u<=6pi so we need to solve sin(u)=-1/2 on [0,6pi] So in the first period [0pi,2pi] we have u=7pi/6 or u=11pi/6 in the second period [2pi,4pi] we have u=2pi+(7pi/6) or u=2pi+(11pi/6) in the third period [4pi,6pi] we have u=4pi+(7pi/6) or u=4pi+(11pi/6) now we need to remember our actual objective solve for x recall we said u IS 3x so in all of these equations replace u with 3x since we made u equal 3x so you have the following 6 equations: 3x=7pi/6 3x=11pi/6 3x=2pi+(7pi/6) 3x=2pi+(11pi/6) 3x=4pi+(7pi/6) 3x=4pi+(11pi/6) now solve for x by dividing each of them by 3.